# Learning Cross-Task Generalities Across Graphs via Task-trees

### v1

@@ Abstract
Foundation models aim to create general, cross-task, and cross-domain machine learning models by pretraining on large-scale datasets to capture shared patterns or concepts (generalities), such as contours, colors, textures, and edges in images, or tokens, words, and sentences in text. However, discovering generalities across graphs remains challenging, which has hindered the development of graph foundation models. To tackle this challenge, in this paper, we propose a novel approach to learn generalities across graphs via task-trees. Specifically, we first define the basic learning instances in graphs as task-trees and assume that the generalities shared across graphs are, at least partially, preserved in the task-trees of the given graphs. To validate the assumption, we first perform a theoretical analysis of task-trees in terms of stability, transferability, and generalization. We find that if a graph neural network (GNN) model is pretrained on diverse task-trees through a reconstruction task, it can learn sufficient transferable knowledge for downstream tasks using an appropriate set of fine-tuning samples. To empirically validate the assumption, we further instantiate the theorems by developing a cross-task, cross-domain graph foundation model named Graph generality Identifier on task-Trees (GIT). The extensive experiments over 30 graphs from five domains demonstrate the effectiveness of GIT in fine-tuning, in-context learning, and zero-shot learning scenarios. Particularly, the general GIT model pretrained on large-scale datasets can be quickly adapted to specific domains, matching or even surpassing expert models designed for those domains. Our data and code are available at https://github.com/Zehong-Wang/GIT.
@@ 1 Introduction
Foundation models have gained prominence in the era of artificial general intelligence as a general-purpose, cross-task, and cross-domain machine learning approach. These models are exemplified by large language models (LLMs) for text (Achiam et al., 2023; Touvron et al., 2023) and large vision models (LVMs) for images (He et al., 2022; Yuan et al., 2021). Pretrained on large-scale datasets, they capture generalizable and transferable knowledge, including contours, colors, textures, and edges in images, as well as tokens, words, and sentences in text. These patterns and concepts represent modality-specific generalities, allowing the models to adapt rapidly to new downstream tasks via in-context learning (Xie et al., 2022) or zero-shot learning (Wei et al., 2021).
@@
Despite the success of foundation models across various modalities, their development of graph-structured data remains in its infancy (Liu et al., 2024). This is primarily due to the high degree of variability among graph-structured datasets (Mao et al., 2024), making it challenging to identify shared generalities across graphs. In particular, graphs from different domains typically represent distinct phenomena; for example, social networks capture relationships between people (Freeman, 2004), while molecular networks depict the structures of molecules (Zeng et al., 2022). These distinctions are evident not only in the differences between feature and label spaces (Huang et al., 2023; Liu et al., 2024) but also in structural heterogeneity (Qiu et al., 2020; Wang et al., 2024d). Additionally, different graph-related tasks often do not share basic learning instances—such as nodes in node-level tasks and entire graphs in graph-level tasks—making it difficult to use a single model for diverse tasks (Mao et al., 2024). The above challenges make it extremely difficult to develop a graph foundation model (GFM) that can identify generalities applicable across graphs.
@@
Is it possible to identify generalities across graphs? Despite the challenges, some researchers have attempted to uncover shared generalities across graphs, which can be mainly categorized into two groups. (1) One line of works draws on graph theory. In particular, they use the concept of graphon (Ruiz et al., 2020) to describe transferable patterns across graphs. If two graphs are generated from the same graphon, they are expected to share key topological properties, which can result in high transferability between them (Ruiz et al., 2020; Cao et al., 2023). However, the strong assumptions underlying graphon theory are often unrealistic for real-world graphs (Levie et al., 2021), and even when the assumptions hold, finding the graphon shared from a large set of graphs presents another significant challenge. (2) Another approach involves leveraging substructures that are present across graphs, such as triangles, stars, and $k$-cliques (Wang et al., 2024c; Zhao et al., 2023; Mao et al., 2024). For instance, triangle structures frequently appear in citation networks, social networks, and molecular networks, albeit with varying semantics. Based on this observation, some works (Sun et al., 2023; Liu et al., 2024) sample subgraphs consisting of these substructures (as shown in Figure 1, Left) and then use a graph neural network (GNN) to encode the subgraph-level embeddings for prediction. However, identifying which substructures are shared across distinct graphs remains difficult. Even if useful substructures can be identified, message-passing GNNs are proved to struggle with capturing such basic substructures in learning (sub)graph embeddings (Garg et al., 2020; Esser et al., 2021; Zhang et al., 2024a), limiting the practicality of these methods.
Figure 1: The formulation of task-trees.
@@
Given the limitations of existing approaches, we propose a novel perspective to answer the question by focusing on the learning process of GNNs. In message-passing GNNs (Kipf & Welling, 2017; Hamilton et al., 2017), predictions are made based on the so-called task-relevant nodes in the graph. For node-level tasks, the task-relevant node is the node itself; for edge-level tasks, it includes the start and end nodes of the edge; and for graph-level tasks, all nodes in the graph are task-relevant. Regardless of the task, basic GNN models only aggregate the embeddings of task-relevant nodes for predictions (Srinivasan & Ribeiro, 2020). This aggregation process is analogous to introducing a virtual “task node” that connects all task-relevant nodes, and learn the embedding of the task nodes for prediction. We refer the computation tree surrounding the “task node” as task-tree, as illustrated in Figure 1 (Right). The task-tree formulation offers three distinct advantages. (1) Learnability: the information within task-trees can be fully captured by message-passing GNNs (Gupta et al., 2024), allowing the model to encode diverse structural information across domains. (2) Uniformity: task-trees are applicable to node-, edge-, and graph-level tasks, providing a unified task alignment paradigm across graphs (Sun et al., 2023). (3) Efficiency: encoding task-trees involves learning the embeddings of virtual nodes appended to the original graph, without extra time-consuming operations. Building on these benefits, it is feasible for task-trees to serve as the basic learning instances in graphs, similar to images in vision and sentences in language. Therefore, we propose that task-trees may maintain transferable patterns or concepts shared across graphs from different domains and tasks, leading to the following assumption:
@@
To validate this assumption, we conduct a theoretical analysis to examine the properties of task-trees in terms of stability, transferability, and generalization (Verma & Zhang, 2019; Garg et al., 2020), showing the feasibility of building GFMs based on task-trees. Our key finding is that if a GNN model is pretrained on diverse task-trees using a reconstruction task, it will acquire sufficient transferable knowledge to downstream tasks with a reasonable number of fine-tuning samples. Additionally, the analysis highlights the potential for specializing the pretrained model into specific domains through post-training (Wei et al., 2021) using domain-specific instances.
To validate our theoretical findings, we introduce a cross-task, cross-domain graph foundation model called Graph generality Identifier on task-Trees (GIT). The core idea is to pretrain GIT on task-trees extracted from a wide range of graphs spanning diverse domains and tasks. We evaluate GIT across 32 graphs from 5 distinct domains, using settings such as basic fine-tuning, in-context learning (few-shot learning without fine-tuning), and zero-shot learning. The experimental results demonstrate that pretraining the model on a small set of graphs significantly benefits a wide range of graph tasks, supporting the idea that task-trees indeed capture generalities shared across graphs. We also propose an instruction tuning method to adapt the pretrained general model to specific downstream domains, which enhances performance across all datasets within the domain, matching or even surpassing domain expert models.
2 Preliminary
We begin with a brief introduction to message-passing GNNs and some related concepts, where a comprehensive discussion on related works is presented in Appendix A. Let ${\mathcal{G}}=({\mathcal{V}},{\mathcal{E}})$ represent a graph with node set ${\mathcal{V}}$ and edge set ${\mathcal{E}}$, where each node $v\in{\mathcal{V}}$ is associated with a feature vector $\mathbf{x}\in\mathbb{R}^{d}$. A GNN encoder $\phi$ takes the graph as input and performs message passing to learn node embeddings ${\bm{Z}}=\phi({\mathcal{V}},{\mathcal{E}})$. Specifically, a GNN encoder can be defined as:
@@ 2 Preliminary
where ${\mathcal{N}}_{i}$ denotes the 1-hop neighbors of node $i$, ${\bm{z}}^{(l)}$ represents the node embedding at the $l$-th GNN layer with ${\bm{z}}^{(0)}={\bm{x}}$, and ${\bm{W}}_{1},{\bm{W}}_{2}$ are learnable matrices. The functions $\sigma$, $\rho$, and $g$ are the activation function, aggregation function and update function, respectively. To simplify the analysis, we assume $\rho$ is an averaging operation and $g$ is the identity function. Without loss of generality (WLOG), these functions can be replaced with any permutation-invariant and Lipschitz-continuous functions, respectively, without affecting the subsequent analysis. We now present some preliminary definitions before outlining the theoretical results.
@@
Graph tasks can be roughly categorized into node-level, edge-level, and graph-level tasks, where the basic learning instances are nodes, edges, and entire graphs, respectively. For node classification, the task-relevant node $v^{t}_{i}$ is the target node $v_{i}$ that needs to be classified. In edge classification, the task-relevant nodes are the start and end nodes $\{v^{t}_{i},v^{t}_{j}\}$ of the target edge $e_{ij}$. For graph classification, the task-relevant nodes $\{v^{t}_{i}\}_{i=1}^{|{\mathcal{V}}|}$ include all nodes in the target graph ${\mathcal{G}}$.
Remark 2.2.
For any graph task instance, the prediction relies solely on the embeddings of the corresponding task-relevant nodes. These node embeddings capture the surrounding subtree structures, also known as computation trees.
@@
The learning process of GNNs involves recursively integrating information from the computation trees, progressing from the bottom to the top. Therefore, for a given node $v$, the node embedding produced by an $L$-layer GNN corresponds to the embedding of its computation tree $T_{v}^{L}$. Since the prediction for any graph task depends solely on the embeddings of task-relevant nodes, and these embeddings are determined by their respective computation trees, we can construct a task-tree for each instance—whether it be a node, edge, or graph—as illustrated in Figure 1 (Right).
@@
For any graph-related instance—whether a node, edge, or graph—we have a set of task-relevant nodes $\{v_{1}^{t},...,v_{n}^{t}\}$ and their corresponding $L$-layer computation trees $\{T_{1},...,T_{n}\}$. These computation trees can be reformulated into a larger task-tree $T^{t}$ by introducing an additional task node that connects all task-relevant nodes.
Task-Tree Encoding. We use a straightforward method to encode task-trees. Given a task-tree $T^{t}$, consisting of a virtual task node $v^{t}$ and a set of task-relevant nodes associated with computation trees ${T_{1},...,T_{n}}$, we first apply a GNN $\phi$ to encode each computation tree, obtaining node embeddings ${{\bm{z}}_{1},...,{\bm{z}}_{n}}$. We then aggregate these node embeddings into the task node, with the task node embedding serving as the embedding of the task-tree. Specifically, we use basic averaging for aggregation: ${\bm{z}}^{t}=\phi(T^{t})=\frac{1}{n}\sum_{i=1}^{n}\phi(T_{i})$.
3 Task-Trees Serve As Basic Learning Instances on Graphs
@@
In this section, we theoretically analyze the properties of task-trees to demonstrate their stability, transferability, and generalization in acting as the basic learning instances, thereby validating the Task-Tree Generality Assumption from a theoretical perspective. It is important to note that the theoretical analysis is not to prove the superiority of task-trees over other methods, such as graphon or subgraph. Rather, we aim to show the feasibility of building GFMs based on task-trees.
Figure 2: subtree information examples.
@@
Given two $L$-layer task-trees $T_{t}^{1}$ and $T_{t}^{2}$, with task-relevant nodes $\{v_{1},...,v_{n}\}$ and $\{v_{1},...,v_{m}\}$, respectively. The distance between task-trees is defined as $\Delta:=\|\phi(T_{1}^{t})-\phi(T_{2}^{t})\|$ with the following bound:
$\Delta=\|\phi(T_{1}^{t})-\phi(T_{2}^{t})\|$ $=\|\frac{1}{n}\sum_{i=1}^{n}\phi(T_{i})-\frac{1}{m}\sum_{j=1}^{m}\phi(T_{j})\|\leq\frac{1}{nm}\sum_{i=1}^{n}\sum_{j=1}^{m}\Bigl{(}{\mathcal{C}}_{1}\|{\bm{x}}_{i}^{(0)}-{\bm{x}}_{j}^{(0)}\|$
$+...+{\mathcal{C}}_{1}{\mathcal{C}}_{2}^{L-1}\|{\bm{x}}_{i}^{(L-1)}-{\bm{x}}_{j}^{(L-1)}\|\Bigr{)}\leq 2{\mathcal{B}}_{\bm{x}}\cdot{\mathcal{C}}_{1}\frac{{\mathcal{C}}_{2}^{L}-1}{{\mathcal{C}}_{2}-1},$ (3)
@@
The proof can be found in Appendix D.1. Theorem 3.1 suggests that two task-trees are likely to have similar representations if their subtrees are similar. This theorem highlights the significance of similarity between pairs of subtrees, while downplaying the impact of the number of subtrees (i.e., the width of the task-trees), despite having more subtrees could potentially increase diversity and thus magnify discrepancy. The theorem also implies that increasing the number of GNN layers may lead to a loose bound, which aligns with previous analyses (Garg et al., 2020; Ju et al., 2023).
@@
Following the stability analysis, we now examine the transferability of task-trees in pretraining and fine-tuning scenario. Specifically, assuming a model is pretrained on a task-tree reconstruction task, we aim to quantify how the knowledge acquired during pretraining can be transferred to downstream tasks. The pretraining objective is defined as ${\mathcal{L}}_{\mathcal{P}}(g\circ\phi):=\mathbb{E}_{(\hat{T},T)\sim{\mathcal{P}}}\|g(\phi(\hat{T}))-\phi(T)\|^{2}$, where ${\mathcal{P}}$ represents the task-tree distribution used for pretraining, $\phi\in\Phi$ and $g\in G$ are the GNN encoder and the reconstruction head, respectively. $T$ denotes the task-tree and $\hat{T}$ is the corrupted version of $T$, generated using arbitrary augmentations. Note that the reconstruction head $g$ is used only during pretraining and is discarded during fine-tuning. Then, we define the risk on downstream task as ${\mathcal{R}}_{\mathcal{T}}(f\circ\phi):=\mathbb{E}_{(T,y)\sim{\mathcal{T}}}\kappa(f(\phi(T)),y)$, where $f\in{\mathcal{F}}$ is a linear head for predictions, ${\mathcal{T}}$ represents the downstream task distribution with task-tree $T$ and label $y$, and $\kappa$ denotes the loss function.
@@
$\min_{f\in{\mathcal{F}}}{\mathcal{R}}_{\mathcal{T}}(f\circ\phi)-\min_{f^{\prime}\in{\mathcal{F}}}{\mathcal{R}}_{\mathcal{T}}(f^{\prime}\circ\phi^{\prime})\leq{\mathcal{C}}_{\delta}\Bigl{(}\min_{g\in G}{\mathcal{L}}_{\mathcal{P}}(g\circ\phi)-\min_{g^{\prime}\in G}{\mathcal{L}}_{\mathcal{P}}(g^{\prime}\circ\phi^{\prime})\Bigr{)}^{\delta},$ (4)
@@
${\mathcal{R}}_{\mathcal{T}}(f\circ\phi)\leq\min_{f^{\prime}\in{\mathcal{F}}}{\mathcal{R}}_{\mathcal{T}}(f^{\prime}\circ\phi^{*})+$ ${\mathcal{C}}_{\delta}\Bigl{(}{\mathcal{E}}_{\mathcal{P}}(g,\phi)\Bigr{)}^{\delta}+\frac{4{\mathcal{C}}_{1}}{n}\sqrt{\sum_{i=1}^{n}\Bigl{\|}\phi(T_{i})\Bigr{\|}^{2}}$
$+$ $2{\mathcal{C}}_{2}\Bigl{(}\sum_{x\in{\mathcal{X}}_{\phi}}\Bigl{\|}{\mathcal{T}}_{\phi}(x)-{\mathcal{P}}_{\phi}(x)\Bigr{\|}+2\sqrt{\frac{\log(1/v)}{n}}\Bigr{)},$ (5)

### v2

# Towards Graph Foundation Models: Learning Generalities Across Graphs via Task-Trees

@@ Abstract
Foundation models aim to create general, cross-task, and cross-domain machine learning models by pretraining on large-scale datasets to capture shared patterns or concepts, such as contours, colors, textures, and edges in images, or tokens, words, and sentences in text. However, identifying generalities across graph-structured data remains a significant challenge, as different graph-based tasks necessitate distinct inductive biases, thereby impeding the development of graph foundation models. To address this challenge, we introduce a novel approach for learning cross-task generalities in graphs. Specifically, we propose task-trees as basic learning instances to align task spaces (node, link, graph) on graphs. Then, we conduct a theoretical analysis to examine their stability, transferability, and generalization. Our findings indicate that when a graph neural network (GNN) is pretrained on diverse task-trees using a reconstruction objective, it acquires transferable knowledge, enabling effective adaptation to downstream tasks with an appropriate set of fine-tuning samples. To empirically validate this approach, we develop a pretrained graph model based on task-trees, termed Graph Generality Identifier on Task-Trees (GIT). Extensive experiments demonstrate that a single pretrained GIT model can be effectively adapted to over 30 different graphs across five domains via fine-tuning, in-context learning, or zero-shot learning. Our data and code are available at https://github.com/Zehong-Wang/GIT.
@@ 1 Introduction
Foundation models have gained prominence in the era of artificial general intelligence as a general-purpose, cross-task, and cross-domain machine learning approach. These models are exemplified by large language models (LLMs) for text (Achiam et al., 2023; Touvron et al., 2023) and large vision models (LVMs) for images (He et al., 2022; Yuan et al., 2021). Pretrained on large-scale datasets, they capture generalizable and transferable knowledge, including contours, colors, textures, and edges in images, as well as tokens, words, and sentences in text. These patterns represent modality-specific generalities, allowing the models to adapt to new downstream tasks via in-context learning (Xie et al., 2022) or zero-shot learning (Wei et al., 2021).
@@
Despite the success of foundation models across various modalities, their development of graph-structured data remains in its infancy (Liu et al., 2024). This is primarily due to the high degree of variability among graph-structured datasets (Mao et al., 2024), making it challenging to identify shared generalities across graphs. In particular, graphs from different domains typically represent distinct phenomena; for example, social networks capture relationships between people (Freeman, 2004), while molecular networks depict the structures of molecules (Zeng et al., 2022). These distinctions are evident not only in feature heterogeneity (Wang et al., 2024b) but also in structural heterogeneity (Qiu et al., 2020; Wang et al., 2024c). More importantly, different graph-related tasks often do not share basic learning instances—such as nodes in node-level tasks and entire graphs in graph-level tasks—making it difficult to use a single model for diverse tasks (Wang et al., 2024b). The above challenges make it difficult to develop a foundation model that can identify generalities applicable across graphs. We focus on the last problem, i.e., task heterogeneity.
@@
Is it possible to identify cross-task generalities across graphs? Despite the challenges, some researchers have attempted to uncover shared generalities across graphs, which can be mainly categorized into two groups. (1) One line of works draws on graph theory. In particular, they use the concept of graphon (Ruiz et al., 2020) to describe transferable patterns across graphs. If two graphs are generated from the same graphon, they are expected to share similar topological properties, which can result in high transferability between them (Ruiz et al., 2020; Cao et al., 2023). However, the strong assumptions underlying graphon theory are often unrealistic for real-world graphs (Levie et al., 2021), and even when the assumptions hold, finding the graphon shared from a large set of graphs presents another significant challenge. (2) Another approach involves leveraging substructures that are present across graphs, such as triangles, stars, and $k$-cliques (Zhao et al., 2023; Mao et al., 2024). For instance, triangle structures frequently appear in citation networks, social networks, and molecular networks, albeit with varying semantics. Based on this observation, some works (Sun et al., 2023; Liu et al., 2024) sample subgraphs consisting of substructures and then use a graph neural network (GNN) to encode the subgraph-level embeddings for prediction. However, message-passing GNNs are proved to struggle with capturing such basic substructures in learning (sub)graph embeddings (Garg et al., 2020; Esser et al., 2021; Zhang et al., 2024a), limiting the practicality of these methods.
@@
Given the limitations of existing approaches, we propose a novel perspective to answer the question by focusing on the learning process of GNNs. In message-passing GNNs (Kipf & Welling, 2017; Hamilton et al., 2017), predictions are made based on the so-called task-relevant nodes in the graph. For node-level tasks, the task-relevant node is the node itself; for edge-level tasks, it includes the start and end nodes of the edge; and for graph-level tasks, all nodes in the graph are task-relevant. Regardless of the task, basic GNN models only aggregate the embeddings of task-relevant nodes for predictions (Srinivasan & Ribeiro, 2020). This aggregation process is analogous to introducing a virtual “task node” that connects all task-relevant nodes, and learn the embedding of the task nodes for prediction. We refer the computation tree surrounding the “task node” as task-tree, as illustrated in Figure 1. The task-tree formulation offers three distinct advantages. (1) Learnability: the information within tree structures can be fully captured by message-passing GNNs (Gupta et al., 2024). (2) Uniformity: task-trees are applicable to node-, edge-, and graph-level tasks, preventing the task heterogeneity on graphs. (3) Efficiency: encoding task-trees technically equals to learning the embeddings of virtual task nodes appended to the original graph, without the need of extracting subgraphs. Building on these benefits, task-trees are able to serve as the basic learning instances in graphs, similar to images in vision and sentences in language. Therefore, task-trees may maintain transferable patterns or concepts shared across different graph tasks, leading to the following assumption:
@@
To validate this assumption, we conduct a theoretical analysis to examine the properties of task-trees in terms of stability, transferability, and generalization. Our key finding is that if a GNN model is pretrained on diverse task-trees using a reconstruction task, it will acquire sufficient transferable knowledge to downstream tasks with a reasonable number of fine-tuning samples. Additionally, the analysis highlights the potential for specializing the pretrained model into specific domains through post-training (Wei et al., 2021) using domain-specific instances.
To empirically verify our theoretical findings, we propose Graph Generality Identifier on Task-Trees (GIT), which is pretrained on task-trees extracted from a wide range of graphs spanning diverse domains and tasks. We evaluate GIT across 32 graphs from 5 distinct domains, via fine-tuning, in-context learning (few-shot learning without fine-tuning), and zero-shot learning. The experimental results demonstrate that pretraining the model on a small set of graphs significantly benefits a wide range of graph tasks, supporting the idea that task-trees indeed capture generalities shared across graph-based tasks. We also propose an instruction tuning method to adapt the pretrained general model to specific downstream domains, which enhances performance across all datasets within the domain, matching or even surpassing domain expert models. We summarize our contributions as follows:
  • We introduce task-trees as basic learning instances for aligning diverse graph-based tasks, demonstrating their superiority over existing instances such as subgraphs.
  • We provide a comprehensive theoretical analysis of the effectiveness of task-trees as the learning instances for task alignment. To the best of our knowledge, this is the first theoretical framework addressing task heterogeneity in graph learning.
  • Leveraging task-trees, we propose GIT, a pretrained model that captures generalizable knowledge across graphs and can be adapted to specialized domains.
  • Extensive experiments conducted on 32 diverse graphs validate the effectiveness of GIT across multiple learning paradigms, including fine-tuning, in-context learning, and zero-shot learning.
2 Task-Trees: Rethinking Basic Learning Tokens in Graph Task Alignment
2.1 Preliminary
We begin with a brief introduction to message-passing GNNs and some related concepts. Let ${\mathcal{G}}=({\mathcal{V}},{\mathcal{E}})$ represent a graph with node set ${\mathcal{V}}$ and edge set ${\mathcal{E}}$, where each node $v\in{\mathcal{V}}$ is associated with a feature vector $\mathbf{x}\in\mathbb{R}^{d}$. A GNN encoder $\phi$ takes the graph as input and performs message passing to learn node embeddings ${\bm{Z}}=\phi({\mathcal{V}},{\mathcal{E}})$. Specifically, a GNN encoder can be defined as:
@@ 2.1 Preliminary
where ${\mathcal{N}}_{i}$ denotes the 1-hop neighbors of node $i$, ${\bm{z}}^{(l)}$ represents the node embedding at the $l$-th GNN layer with ${\bm{z}}^{(0)}={\bm{x}}$, and ${\bm{W}}_{1},{\bm{W}}_{2}$ are learnable matrices. The functions $\sigma$, $\rho$, and $g$ are the activation function, aggregation function and update function, respectively. To simplify the analysis, we assume $\rho$ is an averaging operation and $g$ is the identity function. Without loss of generality (WLOG), these functions can be replaced with any permutation-invariant and Lipschitz-continuous functions, respectively, without affecting the analysis in the paper.
@@
Graph tasks can be roughly categorized into node-level, edge-level, and graph-level tasks, where the basic learning instances are nodes, edges, and entire graphs, respectively. For node classification, the task-relevant node $v^{t}_{i}$ is the node to be classified. In edge classification, the task-relevant nodes are the start and end nodes $\{v^{t}_{i},v^{t}_{j}\}$ of the target edge $e_{ij}$. For graph classification, the task-relevant nodes $\{v^{t}_{i}\}_{i=1}^{|{\mathcal{V}}|}$ include all nodes in the target graph ${\mathcal{G}}$.
For any graph task instance, the prediction relies solely on the embeddings of the corresponding task-relevant nodes. These node embeddings capture the surrounding subtree structures, which are also known as computation trees.
@@
2.2 Task-Tree Construction and Encoding
The learning process of GNNs involves recursively integrating information from the computation trees, progressing from the bottom to the top. Therefore, for a given node $v$, the node embedding produced by an $L$-layer GNN corresponds to the embedding of its computation tree $T_{v}^{L}$. Since the prediction for any graph task depends solely on the embeddings of task-relevant nodes, and these embeddings are determined by their respective computation trees, we can construct a task-tree for each instance—whether it be a node, edge, or graph—as illustrated in Figure 1.
@@ 2.2 Task-Tree Construction and Encoding
For any graph instance—whether a node, edge, or graph—we have a set of task-relevant nodes $\{v_{1}^{t},...,v_{n}^{t}\}$ and their corresponding $L$-layer computation trees $\{T_{1},...,T_{n}\}$. These computation trees can be reformulated into a larger task-tree $T^{t}$ by introducing a virtual node that connects all task-relevant nodes.
We adopt a straightforward approach to encode task-trees. Given a task-tree $T^{t}$, which consists of a virtual task node $v^{t}$ and a set of task-relevant nodes $\{v_{1}^{t},\dots,v_{n}^{t}\}$, we employ a MEAN aggregator for encoding:
${\bm{z}}^{t}=\phi(T^{t})=\frac{1}{n}\sum_{i=1}^{n}\phi(T_{i}),$ (2)
where $T_{i}$ represents the computation tree associated with the task-relevant node $v_{i}^{t}$, and $\phi$ denotes the GNN encoder.
Figure 1: The formulation of task-trees.
Table 1: Comparison of task alignment methods on graphs.
Without Task Alignment Alignment via Subgraph Tasks Alignment via Task-Tree Tasks (Ours)
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Input Original graphs Extracted subgraphs Original graphs with task nodes
Pre-processing N/A Extract subgraphs for each learning instance (node/link/graph) Append task nodes to the original graphs, connecting each to task-relevant nodes
Encoding Apply GNNs to learn node embeddings Apply GNNs on induced subgraphs to obtain subgraph embeddings Apply GNNs directly on original graphs to learn task node embeddings
Prediction Use node embeddings to derive node/link/graph representations Use subgraph embeddings as representations for corresponding instances (node/link/graph) Use task node (task-tree) embeddings as instance representations (node/link/graph) for predictions
Learnability - GNNs struggle to capture fundamental substructures preserved in (sub)graphs GNNs inherently learn from tree structures, such as computation trees
Efficiency - Storing and encoding subgraphs introduce additional computational overhead Encoding task-trees involves encoding augmented virtual nodes in the original graphs with minimal computational cost
Theoretical Basis - Lack of a well-established theoretical foundation Supported by a rigorous theoretical framework
2.3 Comparison to Existing Works
Unlike our proposed task-trees, several existing approaches (Qiu et al., 2020; Sun et al., 2023; Huang et al., 2023; Liu et al., 2024; He & Hooi, 2024) utilize $k$-hop subgraphs extracted from graphs as the basic learning instances. For instance, in node classification, ego-graphs are constructed around each node, where the label of the central node is assigned to the induced subgraph, effectively reformulating node classification as a subgraph classification task. A similar transformation can be applied to edge-level and graph-level tasks by converting them into subgraph-level learning problems. This method involves: (1) extracting ego-graphs centered around task-relevant nodes and (2) applying GNNs to learn graph-level embeddings for classification. However, this subgraph extraction process incurs substantial computational overhead, increasing both time and memory requirements due to the necessity of storing and processing induced subgraphs. Moreover, information within these subgraphs is not always effectively captured by message-passing GNNs, as GNNs may struggle to learn essential substructures preserved in graphs (Garg et al., 2020; Chen et al., 2020; Zhang et al., 2024a), thereby limiting the efficacy of subgraphs as learning instances.
In contrast, our proposed task-trees offer both greater efficiency and improved learnability. Specifically, encoding task-trees for node, link, or graph-level tasks involves: (1) augmenting the original graph by adding virtual task nodes and connecting them to task-relevant nodes and (2) applying GNNs over the augmented graph to encode the embeddings of these virtual nodes for prediction. For example, in node classification, we first introduce virtual nodes connected to each node in the original graph and subsequently apply GNNs to this augmented structure, allowing virtual node embeddings to be learned for classification. Consequently, our method requires only the addition of nodes and edges to the existing graph, making it significantly more efficient than extracting and storing subgraphs. Furthermore, encoding task-trees is equivalent to directly encoding virtual nodes through message passing, ensuring that task-tree information remains fully learnable by standard GNNs. Empirically, task-trees consistently outperform subgraphs in both effectiveness and efficiency (Section 5.5 and Appendix I). A detailed comparison is presented in Table 1, with additional discussions on related work available in Appendix A.
3 Why Do Task-Trees Serve as Basic Learning Instances?
@@
In this section, we provide a theoretical analysis of task-trees, highlighting their stability, transferability, and generalization properties as basic learning instances. This analysis serves to validate the Task-Tree Generality Assumption from a theoretical standpoint. Importantly, our objective is not to claim the superiority of task-trees over other approaches, such as subgraphs. Rather, we establish the feasibility of learning cross-task generalities based on task-trees.
@@
Given two $L$-layer task-trees $T_{t}^{1}$ and $T_{t}^{2}$, with task-relevant nodes $\{v_{1},...,v_{n}\}$ and $\{v_{1},...,v_{m}\}$, respectively. The distance between task-trees is defined as $\Delta:=\|\phi(T_{1}^{t})-\phi(T_{2}^{t})\|$ with
$\Delta$ $=\|\phi(T_{1}^{t})-\phi(T_{2}^{t})\|=\|\frac{1}{n}\sum_{i=1}^{n}\phi(T_{i})-\frac{1}{m}\sum_{j=1}^{m}\phi(T_{j})\|$
$\leq\frac{1}{nm}\sum_{i=1}^{n}\sum_{j=1}^{m}\Bigl{(}{\mathcal{C}}_{1}\|{\bm{x}}_{i}^{(0)}-{\bm{x}}_{j}^{(0)}\|+...$ (4)
$+{\mathcal{C}}_{1}{\mathcal{C}}_{2}^{L-1}\|{\bm{x}}_{i}^{(L-1)}-{\bm{x}}_{j}^{(L-1)}\|\Bigr{)}\leq 2{\mathcal{B}}_{\bm{x}}\cdot{\mathcal{C}}_{1}\frac{{\mathcal{C}}_{2}^{L}-1}{{\mathcal{C}}_{2}-1},$
@@
Theorem 3.1 (proved in Appendix D.1) suggests that two task-trees are likely to have similar representations if their subtrees are similar. This theorem highlights the significance of similarity between pairs of subtrees, while downplaying the impact of the number of subtrees (i.e., the width of the task-trees), despite having more subtrees could potentially increase diversity and thus magnify discrepancy. The theorem also implies that increasing the number of GNN layers may lead to a loose bound, which aligns with previous analyses (Garg et al., 2020; Ju et al., 2023).
@@
We now examine the transferability of task-trees. Specifically, assuming a model is pretrained on a task-tree reconstruction task [^1] The scope of reconstruction task is large. We consider contrastive learning is also a kind of reconstruction., we aim to quantify how the knowledge acquired during pretraining can be transferred to downstream tasks. The pretraining objective is defined as ${\mathcal{L}}_{\mathcal{P}}(g\circ\phi):=\mathbb{E}_{(\hat{T},T)\sim{\mathcal{P}}}\|g(\phi(\hat{T}))-\phi(T)\|^{2}$, where ${\mathcal{P}}$ represents the task-tree distribution used for pretraining, $\phi\in\Phi$ and $g\in G$ are the GNN encoder and the reconstruction head, respectively. $T$ denotes the task-tree and $\hat{T}$ is the corrupted version of $T$, generated using arbitrary augmentations. Note that the reconstruction head $g$ is used only during pretraining and is discarded during fine-tuning. Then, we define the risk on downstream task as ${\mathcal{R}}_{\mathcal{T}}(f\circ\phi):=\mathbb{E}_{(T,y)\sim{\mathcal{T}}}\kappa(f(\phi(T)),y)$, where $f\in{\mathcal{F}}$ is a linear head for predictions, ${\mathcal{T}}$ represents the downstream task distribution with task-tree $T$ and label $y$, and $\kappa$ denotes the loss function.
@@
$\leq{\mathcal{C}}_{\delta}\Bigl{(}\min_{g\in G}{\mathcal{L}}_{\mathcal{P}}(g\circ\phi)-\min_{g^{\prime}\in G}{\mathcal{L}}_{\mathcal{P}}(g^{\prime}\circ\phi^{\prime})\Bigr{)}^{\delta},$ (5)
@@
${\mathcal{R}}_{\mathcal{T}}$ $(f\circ\phi)\leq\min_{f^{\prime}\in{\mathcal{F}}}{\mathcal{R}}_{\mathcal{T}}(f^{\prime}\circ\phi^{*})$
$+2{\mathcal{C}}_{2}\Bigl{(}\sum_{x\in{\mathcal{X}}_{\phi}}\Bigl{\|}{\mathcal{T}}_{\phi}(x)-{\mathcal{P}}_{\phi}(x)\Bigr{\|}+2\sqrt{\frac{\log(1/v)}{n}}\Bigr{)}$
$+{\mathcal{C}}_{\delta}\Bigl{(}{\mathcal{E}}_{\mathcal{P}}(g,\phi)\Bigr{)}^{\delta}+\frac{4{\mathcal{C}}_{1}}{n}\sqrt{\sum_{i=1}^{n}\Bigl{\|}\phi(T_{i})\Bigr{\|}^{2}},$ (6)